# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Problem-Space Evasion Adversarial Attacks on Flow-based ML-NIDS

### v3

@@ 3.2.4 PS4[$f_{i}$]: Features that attackers cannot access
C3: the feature $f_{i}$ is not a backward or interflow feature. In other words, attackers can access $f_{i}$.
C4: the feature $f_{i}$ is a backward or interflow feature; however, it is highly correlated with a forward feature. In other words, attackers can modify $f_{i}$ in an indirect way.
C5: the feature $f_{i}$ is a backward or interflow feature; however, it is correlated with multiple forward features. In other words, attackers can modify $f_{i}$ indirectly, but it will be challenging for them as it is correlated with multiple features.
Otherwise (if none of C3, C4, or C5 apply): the feature $f_{i}$ is a backward or interflow feature and it is not correlated with any forward feature. In other words, attackers cannot access $f_{i}$.
@@
3.2.5 PS5[$f_{i}$]: Flow-wide Correlation
@@
C6: $f_{i}$ is a feature that requires modifying the entire flow of packets (forward, backward, or both), such as mean or standard deviation features.
@@ 3.2.6 PSTotal[$f_{i}$]
3.3 Ranges of PS Fields and Thresholds Calibration
@@
As previously discussed, the total perturb-ability score (PSTotal[$f_{i}$]) for each feature $f_{i}$ is computed as the geometric mean of the five individual PS fields, ensuring a balanced assessment across different conditions. This approach was adopted to account for the fact that certain feature properties can entirely negate their perturb-ability, such as when the feature is inaccessible to the attacker. These features are given a PS of 0. On the other hand, other conditions may only reduce the feature’s PS without fully eliminating its susceptibility. For instance, features that require altering an entire packet flow—whether forward, backward, or both—such as mean or standard deviation of packet properties, tend to decrease the overall PS score without completely nullifying it.
Figure 7: Using PS as a Potential Defense against Practical Problem-Space Evasion Adversarial Attacks
@@
While certain thresholds, such as 255 in $\text{PS}_{2}[f_{i}]$ and 10 in $\text{PS}_{3}[f_{i}]$, might appear arbitrary at first glance, in the previous section, we provided a detailed rationale for these choices to the best of our ability. These thresholds were selected based on empirical evidence by calibrating PS thresholds against the datasets we used and domain-specific considerations, aiming to balance the practical limitations and the adversarial potential of each feature. However, it is important to recognize that these thresholds, along with the underlying functions of the PS fields, are not set in stone.
As with many design threshold decisions, these parameters and thresholds can—and often should—be calibrated, adapted and fine-tuned by machine learning engineers, domain experts, and practitioners prior to deployment. The flexibility to adjust such thresholds ensures that the perturb-ability scoring system remains both robust and adaptable to varying real-world scenarios, where the specific characteristics of the network and potential threat models may differ. This adaptability mirrors the iterative nature of system optimization in machine learning pipelines, where parameter tuning plays a pivotal role in enhancing overall performance and security. By allowing these modifications, the PS framework is better equipped to handle the dynamic and evolving landscape of adversarial attacks in network security systems.
3.4 PS as a Defense
@@
The development of a Machine Learning-based NIDS (ML-NIDS) begins with network packet data undergoing a rigorous sequence of pre-processing steps. These steps include feature extraction, data cleansing, outlier exclusion, standardization, and one-hot encoding. Once pre-processing is complete, the resulting feature vectors are divided into balanced training sets, as well as validation and testing sets that maintain the original data distribution.
During the training phase, various machine learning models, such as Artificial Neural Networks (ANN), Convolutional Neural Networks (CNN), and Support Vector Machines (SVM), are evaluated for performance. Hyperparameter tuning and the selection of training evaluation metrics further optimize these models to ensure robust detection capabilities. In the testing phase, model performance is assessed through accuracy and F1 score metrics, with results compared against ground truth labels to measure the system’s robustness and reliability.
@@
A distinguishing aspect of our approach is the exclusion of features with high perturb-ability scores during the feature selection process. By doing so, attackers encounter no or very few perturb-able features in problem-space, significantly reducing the attack surface and making it significantly more difficult for adversaries to exploit the system. This method ensures that the features retained for training and classification are inherently resistant to adversarial manipulations. While this may require rethinking traditional feature selection methods, the potential benefits in preventing evasion attempts are substantial. This simple, efficient solution utilizes NIDS domain constraints as a defense with minimal computational overhead.
4 Experimental Environment
@@
4.1 Datasets
For our experiments, we used 2 public NIDS datasets; (1) UNSW-NB15 [29]: A widely-used NIDS dataset with 2.5M records (2.2M normal, 0.32M attacks) across 9 attack categories. It includes 49 features and represents realistic network traffic and diverse attack vectors. (2) CSE-CIC-IDS2018 (Improved) [28]: A dataset with various attacks (e.g., Brute Force, DoS, DDoS) and benign traffic. We use the improved version by Liu et al. [28], which resolves issues in the original dataset [33] related to feature generation and labeling, ensuring better reliability for research.
4.2 Features’ Analysis and Correlations
@@ 4.3 Pre-processing
Table 1: The number and percentage of features in every perturb-ability class, based on our proposed PS, where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class
@@
4.4 Machine and Deep Learning Models
Table 2: UNSW-NB15 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
@@ 4.4 Machine and Deep Learning Models
Table 3: CSE-CIC-IDS2018 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
@@
5 Results
5.1 Evaluation Plan for PS
In this subsection, we outline our evaluation plan for the PS metric, focusing on several research questions (RQs) that guide our investigation into the effectiveness and applicability of PS in enhancing the robustness of ML-NIDS;
RQ1: How is the distribution of features characterized using PS? (section 5.2)
RQ2: To which PS class do the ML-NIDS features belong? (section 5.2)
RQ3: What cutoff is used to differentiate between PS classes? and why was it used? (section 5.2)
RQ4: Can ML-NIDS models utilizing only green (low perturb-ability) or a combination of green and yellow (low and medium perturb-ability) features perform as effectively as models that include all features? (section 5.3)
RQ5: Does PS effectively capture the easily perturb-able features used in problem-space adversarial attacks against NIDS research? (section 5.4)
@@ 5.1 Evaluation Plan for PS
5.2 Classification of the Features Using PS (RQ1,2,3)
Table 4: The performance of an ANN/Random Forest (RF)/SVM/CNN-based NIDS
@@ 5.2 Classification of the Features Using PS (RQ1,2,3)
with only the low Pert. SVM 0.9879 0.9129 0.9996 0.9543 0.9999 0.9984 0.9994 0.9989
@@
  (c) Performance of ML-NIDS ANN 0.9880 0.9130 0.9999 0.9545 1.0000 0.9996 0.9998 0.9997
selected during features selection CNN 0.9882 0.9145 0.9997 0.9552 1.0000 0.9994 1.0000 0.9997
Table 1 presents the distribution of features from the two utilized datasets (UNSW-NB15 and the improved CSE-CIC-IDS2018) across different perturb-ability classes, based on our proposed perturb-ability Score (PS). Features classified as low perturb-ability (green) have a PS of 0, indicating robustness to adversarial manipulation. High perturb-ability features (red) are characterized by a PS greater than or equal to 0.87, marking them as highly perturb-able. We selected the threshold of 0.87 to define highly perturb-able features, as discussed earlier. This threshold is derived from the geometric mean of the five PS components, ensuring that at most one PS field has a value as low as 0.5. In our PS evaluation criteria, a score of 0.5 signifies a condition that tends to reduce the overall PS score, without entirely nullifying it. Using the 0.87 threshold ensures that a feature with at most one such condition is still classified within the high perturb-ability class, thereby reinforcing the rigor of our feature selection criteria. The remaining features, which are neither in the high nor low perturb-ability classes, fall into the medium perturb-ability class (yellow). Table 1 also shows the percentages of these features across the perturb-ability classes, as defined by our PS. Approximately 35% of the features in both datasets belong to the high perturb-ability class (red). Fig. 10 depicts the histogram of PS values for each dataset, with the Y-axis representing the number of features and the X-axis representing the PS score. The distribution of PS scores is illustrated, along with the cutoffs between the classes. Table LABEL:fig:NB15PS provides a detailed classification of the UNSW-NB15 dataset features according to their perturb-ability scores. Features such as ‘dstip‘ and ‘service‘ are categorized as low perturb-ability (green), implying their robustness against adversarial manipulations. In contrast, features like ‘dur‘ and ‘stime‘ are identified as high perturb-ability (red), indicating their susceptibility to adversarial attacks. Table 3 presents the feature classification for the improved CSE-CIC-IDS2018 dataset. Medium perturb-ability features (yellow), such as ‘Flow Bytes/s‘, occupy an intermediate position between these extremes. It is important to note that low perturb-ability (green) features do not imply that an attacker cannot modify them. Instead, modifying these features may disrupt the network or interfere with the malicious functionality of an attack, or the attacker may have limited or no access to them. Tables 6, 7, and 8 in Appendix A show the definitions of the features in the used datasets.
5.3 Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ4)
Table 4 presents the performance of ANN, Random Forest, SVM, and CNN as classifiers in NIDS. The evaluation is conducted on two datasets, with results shown for various feature combinations based on their perturb-ability. The table is structured into three sections: (a) performance with low perturb-ability features (green), (b) performance with low and medium perturb-ability features (green + yellow), and (c) performance with all features (green + yellow + red).
For both datasets, the performance remains consistent across all feature combinations, which is highly promising as it suggests that selecting only low perturb-ability (green) features can still yield excellent model performance. The improved CSE-CIC-IDS2018 dataset demonstrated outstanding performance across all models, metrics, and feature sets. Similarly, the UNSW-NB15 dataset performed very well across all models, metrics, and selected features, although precision was slightly lower (but still above 0.912, which is excellent) compared to the other dataset. However, this marginal difference in precision was unrelated to our perturb-ability-based feature selection. Precision remained consistent across all selected feature sets. Additionally, we ranked the features using Random Forest to select the most important ones, achieving similar precision results. We also experimented with various balancing techniques, such as SMOTE [13], but observed no significant difference in precision. This consistency further confirms that the precision difference is independent of our perturb-ability-based feature selection. In a nutshell, Table 4 shows that using only low perturb-ability (green) features does not create low-performance models; on the contrary, all of these models have an accuracy and F1 score above 0.9.
It is important to note that in all three cases (a, b, and c) of Table 4, thorough pre-processing was conducted, as discussed in Section section 4.3, and feature engineering was applied to remove spurious or faulty features. For example, IP addresses were not treated as numeric values, as this does not generalize well; instead, basic geolocation information was extracted. Time-based features like timestamps were also dropped, as attacks can happen at any time, making this feature non-generalizable. Identification features (e.g., flow ID) were removed for the same reason. Additionally, we applied feature importance ranking techniques to identify the most influential features in the model, ensuring that these attributes provide meaningful and predictive information about the target variable.
@@ 5.3 Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ4)
Table 5: Mapping problem-space evasion adversarial attacks’ traffic morphing techniques to features, the features are colored based on our PS classification.
@@
5.4 Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ5)
To validate our proposed PS, we analyzed problem-space evasion adversarial attacks targeting ML-NIDS [21, 22, 37, 38, 36, 39, 24, 8]. Table 5 presents a detailed mapping of traffic morphing techniques employed in these problem-space adversarial attacks to the corresponding perturbed features within the feature-space of the two datasets utilized in this study. The table outlines the problem-space morphing techniques and maps them to the potentially impacted features. Additionally, it classifies these features according to our perturb-ability Score (PS) system, using a color-coded scheme: low perturb-ability (green), medium perturb-ability (yellow), and high perturb-ability (red).

### v4

@@ III-B4 PS4[$f_{i}$]: Features that attackers cannot access
C4: the feature $f_{i}$ is not a backward or interflow feature. In other words, attackers can access $f_{i}$.
C5: the feature $f_{i}$ is a backward or interflow feature; however, it is highly correlated with a forward feature. In other words, attackers can modify $f_{i}$ in an indirect way.
C6: the feature $f_{i}$ is a backward or interflow feature; however, it is correlated with multiple forward features. In other words, attackers can modify $f_{i}$ indirectly, but it will be challenging for them as it is correlated with multiple features.
Otherwise (if none of C4, C5, or C6 apply): the feature $f_{i}$ is a backward or interflow feature and it is not correlated with any forward feature. In other words, attackers cannot access $f_{i}$.
@@
III-B5 PS5[$f_{i}$]: Flow-wide Correlation
@@
C7: $f_{i}$ is a feature that requires modifying the entire flow of packets (forward, backward, or both), such as mean or standard deviation features.
@@ III-B6 PSTotal[$f_{i}$]
III-C Ranges of PS Fields and Thresholds Calibration
@@
As previously discussed, the total perturb-ability score (PSTotal[$f_{i}$]) for each feature $f_{i}$ is computed as the geometric mean of the five individual PS fields, ensuring a balanced assessment across different conditions. This approach was adopted to account for the fact that certain feature properties can entirely negate their perturb-ability, such as when the feature is inaccessible to the attacker. These features are given a PS of 0. On the other hand, other conditions may only reduce the feature’s PS without fully eliminating its susceptibility. For instance, features that require altering an entire packet flow, whether forward, backward, or both, such as mean or standard deviation of packet properties, tend to decrease the overall PS score without completely nullifying it.
Figure 7: Option A Defense: PS-enabled Feature Selection
@@
While certain thresholds, such as 2 (MinR) and 255 (MaxR) in $\text{PS}_{2}[f_{i}]$, might appear arbitrary at first glance, in the previous section, we provided a detailed rationale for these choices to the best of our ability. However, it is important to recognize that these thresholds, along with the underlying functions of the PS fields, are not set in stone. As with many design threshold decisions, these parameters and thresholds can, and often should, be calibrated, adapted and fine-tuned by machine learning engineers, domain experts, and practitioners prior to deployment. The flexibility to adjust such thresholds ensures that the perturb-ability scoring system remains both robust and adaptable to varying and dynamic real-world scenarios, where the specific characteristics of the network and potential threat models may differ.
IV PS-enabled Defenses
In this section, we introduce two methods where PS enables defenses against practical problem-space evasion adversarial attacks against ML-NIDS.
IV-A Option A: PS-enabled Feature Selection
@@
By the exclusion of features with high perturb-ability scores during the feature selection process during the development of an ML-NIDS, attackers encounter no or very few perturb-able features in problem-space, significantly reducing the attack surface and making it significantly more difficult for adversaries to exploit the system. This method ensures that the features retained for training and classification are inherently resistant to adversarial manipulations. While this may require rethinking traditional feature selection methods, the potential benefits in preventing evasion attempts are substantial. This simple, efficient solution utilizes NIDS domain constraints as a defense with minimal computational overhead.
IV-B Option B: PS-enabled Feature Masking
Unlike our Option A defense, which requires retraining the entire model on a selected subset of features if the model is already running, Option B provides a more efficient approach when the system is already deployed and the cost of re-selecting features and re-training is prohibitively high. Our novel approach allows us to maintain the existing model architecture and dimensionality while masking high perturb-ability features.
Our feature masking approach builds upon the binary feature mask optimization framework introduced by Lorasdagi et al. [26], but with a critical distinction. Their approach is primarily for feature selection during model development, while ours is designed as a defense mechanism for already deployed systems.
We define a mask vector $\mathbf{m}\in{0,1}^{M}$, where $M$ is the number of features in the dataset. The mask is determined by the PS scores:
$m_{i}=\begin{cases}0,&\text{if }\text{PS}_{\text{Total}}[f_{i}]\geq\tau\\ 1,&\text{otherwise}\end{cases}$
where $\tau$ is a threshold that determines which features are considered high perturb-ability. Unlike binary feature mask optimization, which applies the mask through Hadamard product (element-wise multiplication), our approach replaces masked features with neutral values:
$x_{i}^{\prime}=\begin{cases}\nu_{i},&\text{if }m_{i}=0\\ x_{i},&\text{otherwise}\end{cases}$
where $x_{i}$ is the original feature value, $x_{i}^{\prime}$ is the masked feature value, and $\nu_{i}$ is a neutral value for feature $i$ (typically the mean or median of that feature from the training data).
By using neutral values instead of always zeros, our approach aligns with feature amputation techniques, which have been shown to be more effective in maintaining model performance when dealing with missing or corrupted features. This is particularly important in the context of adversarial defense, where we aim to minimize the impact on legitimate traffic classification while maximizing robustness against attacks. Moreover, replacing with neutral values maintains the overall statistical distribution of the data, reducing the risk of introducing artifacts that could affect model performance. The neutral values $\nu_{i}$ could be a single value for all features (for example, 0.5 if the model uses min-max normalization or 0 if the model uses standardization), or it can be a designated value per feature computed from the training data (e.g., mean or median) and stored for use during inference. In our experiments, we tested multiple values of $\nu_{i}$ to validate our work.
For our Option B defense, we test two variants: B1 and B2. Option B1 integrates feature masking during both training and inference, selectively suppressing high-perturbability features while preserving the original model architecture and input dimensionality. This eliminates the need for pipeline restructuring or retraining on reduced feature subsets (as required by Option A), ensuring compatibility with deployed systems. Option B2, in contrast, applies masking exclusively during inference by replacing high-PS features with precomputed neutral values (e.g., training-data means or medians). This inference-only approach incurs negligible computational overhead, requires no retraining, and enables real-time adaptation to evolving adversarial threats by adjusting masked features on-the-fly. While B1 ensures robustness through training-phase integration, B2 prioritizes operational agility, as the system can dynamically adjust which features to mask based on evolving threat landscapes without requiring retraining or modifications to pretrained models. Our PS-enabled feature masking techniques reduce attack surface: By replacing easily perturbable features with neutral values, we effectively remove attackers’ ability to manipulate those features while maintaining the model’s structure.
V Experimental Environment
@@
V-A Datasets
For our experiments, we used 3 public NIDS datasets; (1) UNSW-NB15 [27]: A widely-used NIDS dataset with 2.5M records (2.2M normal, 0.32M attacks) across 9 attack categories. It includes 49 features and represents realistic network traffic and diverse attack vectors. (2) CSE-CIC-IDS2018 (Improved) [28]: A dataset with various attacks (e.g., Brute Force, DoS, DDoS) and benign traffic. We use the improved version by Liu et al. [28], which resolves issues in the original dataset [29] related to feature generation and labeling, ensuring better reliability for research. (3) MCFP [30]: Unlike the two previous datasets which had CSV feature-space formats, MCFP is a problem-space dataset consisting of PCAP files capturing real network traffic. MCFP was created by the Malware Capture Facility Project at Stratosphere Lab and consists of benign and various malware types’ raw network traffic captures in realistic environments, along with their ground-truth information
V-B Features’ Analysis and Correlations
@@ V-C Pre-processing
V-D Machine and Deep Learning Models
VI Research Questions
TABLE I: The number and percentage of features in every perturb-ability class, based on our proposed PS, where green indicates low perturb-ability features class, yellow indicates medium perturb-ability features class, and red indicates high perturb-ability features class
@@ VI Research Questions
TABLE II: UNSW-NB15 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
@@
TABLE III: CSE-CIC-IDS2018 Dataset’s features classified based on our proposed PS, where green indicates a feature with low perturb-ability, yellow indicates a feature with medium perturb-ability, and red indicates a feature with high perturb-ability.
@@
The following are our goals and their associated Research Questions (RQs) for each:
  1. I.
    Knowing the Distribution of Features’ PS Classification
    • RQ1: How are ML-NIDS features characterized and distributed across PS classes, and what criteria determine the classification cutoffs? (section VII)
  2. II.
    Testing the Possibility of Deploying PS-Enabled Defenses
    • RQ2: Can ML-NIDS models utilizing only low perturb-ability (green) features or a combination of low and medium perturb-ability (green and yellow) features perform as effectively as models that include all features? (Option A defense possibility) (section VIII)
    • RQ3: Can ML-NIDS models with high perturb-ability (red) features and/or medium perturb-ability (yellow) features masked perform as effectively as models that maintain all features unmasked? (Option B defenses possibility) (section VIII)
  3. III.
    Validation of PS as an indicator to perturb-able features
    • RQ4: Does PS effectively identify the easily perturb-able features exploited in problem-space adversarial attacks against NIDS documented in research? (section IX)
  4. IV.
    Testing the Effect of PS-Enabled Defenses on Adversarial Attacks
    • RQ5: Can our PS-enabled defenses neutralize adversarial attacks against flow-based ML-NIDS? (section X)
@@
VII Features’ Classification Using PS (RQ1)
Table I presents the distribution of features from the two feature-space datasets (UNSW-NB15 and the improved CSE-CIC-IDS2018) across different perturb-ability classes, based on our proposed perturb-ability Score (PS). Features classified as low perturb-ability (green) have a PS of 0, indicating robustness to adversarial manipulation. High perturb-ability features (red) are characterized by a PS greater than or equal to 0.87, marking them as highly perturb-able. We selected the threshold of 0.87 to define highly perturb-able features, as discussed earlier. This threshold is derived from the geometric mean of the five PS components, ensuring that at most one PS field has a value as low as 0.5. In our PS evaluation criteria, a score of 0.5 signifies a condition that tends to reduce the overall PS score, without entirely nullifying it. Using the 0.87 threshold ensures that a feature with at most one such condition is still classified within the high perturb-ability class, thereby reinforcing the rigor of our feature selection criteria. The remaining features, which are neither in the high nor low perturb-ability classes, fall into the medium perturb-ability class (yellow). Table I also shows the percentages of these features across the perturb-ability classes, as defined by our PS. Approximately 35% of the features in both datasets belong to the high perturb-ability class (red). Fig. 10 depicts the histogram of PS values for each dataset, with the Y-axis representing the number of features and the X-axis representing the PS score. The distribution of PS scores is illustrated, along with the cutoffs between the classes. Table II provides a detailed classification of the UNSW-NB15 dataset features according to their perturb-ability scores. Features such as ‘dstip‘ and ‘service‘ are categorized as low perturb-ability (green), implying their robustness against adversarial manipulations. In contrast, features like ‘dur‘ and ‘stime‘ are identified as high perturb-ability (red), indicating their susceptibility to adversarial attacks. Table III presents the feature classification for the improved CSE-CIC-IDS2018 dataset. Medium perturb-ability features (yellow), such as ‘Flow Bytes/s‘, occupy an intermediate position between these extremes. It is important to note that low perturb-ability (green) features do not imply that an attacker cannot modify them. Instead, modifying these features may disrupt the network or interfere with the malicious functionality of an attack, or the attacker may have limited or no access to them. Tables XIV, XV, and XVI in Appendix C show the definitions of the features in the used datasets.
It is important to note that in all of our experiments, thorough pre-processing was conducted, as discussed in Section section V-C, and feature engineering was applied to remove spurious or faulty features. Time-based features like timestamps were dropped, as attacks can happen at any time, making these features non-generalizable. Identification features (e.g., flow ID), while green, were removed for the same reason.
VIII Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ2,3)
TABLE IV: The Baseline Performance of an ANN/Random Forest (RF)/SVM/CNN-based NIDS
Dataset $\rightarrow$ UNSW-NB15     Improved CSE-CIC-IDS2018
Performance of ML-NIDS ANN 0.9883 0.9158 0.9991 0.9557     1.0000 0.9997 0.9998 0.9997
selected during features selection CNN 0.9888 0.9201 0.9976 0.9573     1.0000 0.9995 0.9999 0.9997
TABLE V: The performance of an ANN/RF/SVM/CNN-based NIDS After PS-enabled Feature Selection (Option A Defense).
@@ VIII Performance of ML-NIDS using various combinations of features based on their perturb-ability (RQ2,3)
with only the low Pert. SVM 0.9879 0.9129 0.9997 0.9543     0.9999 0.9984 0.9994 0.9989
@@
TABLE VI: The performance of an ANN/RF/SVM/CNN-based NIDS after PS-enabled feature masking during training and inference (Option B1 Defense).
Dataset $\rightarrow$ UNSW-NB15     Improved CSE-CIC-IDS2018
(a) Performance of ML-NIDS ANN 0.9879 0.9128 0.9998 0.9543     1.0000 0.9994 0.9999 0.9996
with only the low Pert. SVM 0.9879 0.9130 0.9996 0.9543     0.9999 0.9982 0.9995 0.9988
features (Green) not masked RF 0.9892 0.9225 0.9982 0.9588     1.0000 0.9998 1.0000 0.9999
((Red + Yellow) are masked) CNN 0.9881 0.9142 0.9994 0.9549     1.0000 0.9997 0.9998 0.9998
  (b) Performance of ML-NIDS ANN 0.9879 0.9128 1.0000 0.9544     1.0000 0.9996 1.0000 0.9998
with only the low and med Pert. SVM 0.9879 0.9130 0.9997 0.9543     0.9999 0.9982 0.9998 0.9990
features (Green + Yellow) not masked RF 0.9893 0.9233 0.9981 0.9592     1.0000 0.9998 1.0000 0.9999
(Red are masked) CNN 0.9881 0.9140 0.9998 0.9550     1.0000 0.9997 0.9998 0.9998
TABLE VII: The performance of an ANN/RF/SVM/CNN-based NIDS after PS-enabled feature masking during inference (Option B2 Defense).
Dataset $\rightarrow$ UNSW-NB15     Improved CSE-CIC-IDS2018
(a) Performance of ML-NIDS ANN 0.9862 0.9018 0.9999 0.9483     0.9980 0.9684 0.9999 0.9839
with only the low Pert. SVM 0.9872 0.9079 0.9999 0.9517     0.9988 0.9806 0.9990 0.9897
features (Green) not masked RF 0.9841 0.8883 1.0000 0.9408     0.9971 0.9998 0.9513 0.9749
((Red + Yellow) are masked) CNN 0.9876 0.9110 0.9998 0.9533     0.9977 0.9635 0.9999 0.9814
  (b) Performance of ML-NIDS ANN 0.9862 0.9018 0.9999 0.9483     0.9987 0.9793 0.9998 0.9895
with only the low and med Pert. SVM 0.9880 0.9133 0.9995 0.9545     0.9985 0.9753 0.9996 0.9873
features (Green + Yellow) not masked RF 0.9868 0.9054 1.0000 0.9504     0.9999 0.9999 0.9985 0.9992
(Red are masked) CNN 0.9879 0.9129 0.9993 0.9541     0.9982 0.9717 1.0000 0.9856
@@
TABLE VIII: Mapping problem-space evasion adversarial attacks’ traffic morphing techniques to features, the features are colored based on our PS classification.
@@
To evaluate the validity of our PS-enabled defenses, we conducted extensive experiments using four different machine learning models (ANN, SVM, Random Forest (RF), and CNN) on two feature-space datasets (UNSW-NB15 [27] and Improved CSE-CIC-IDS2018 [28]). These tests aim to check if selecting non-perturb-able features or masking all but non-perturb-able features will significantly reduce the performance of the models during our PS-enabled defenses.
Table IV shows the baseline performance of the ML-NIDS models using all features (Green + Yellow + Red). This serves as our reference point for comparison. As we can observe, all models achieve high performance across both datasets, with accuracy, precision, recall, and F1 scores consistently above 0.99 for the Improved CSE-CIC-IDS2018 dataset and above 0.95 for the UNSW-NB15 dataset.
To address RQ2, we evaluated the performance of ML-NIDS models using only low perturb-ability (Green) features and a combination of low and medium perturb-ability (Green + Yellow) features. The results are presented in Table V. Remarkably, we observe that the performance of the models using only Green features (Table Va) or Green + Yellow features (Table Vb) is nearly identical to the baseline performance using all features. For instance, the ANN model on the UNSW-NB15 dataset maintains an accuracy of 0.9879 and an F1 score of 0.9544 when using only Green features, compared to 0.9883 and 0.9557 respectively when using all features. This trend is consistent across all models and both datasets, demonstrating that ML-NIDS models can perform effectively using only low or low-medium perturb-ability features. Thus, based on these results, implementing PS-enabled option A defense with only the green features selected may be preferable as it further reduces the attack surface without model performance degradation.
To address RQ3, we evaluated PS-enabled feature masking under two scenarios: masking during both training and inference (Table VI) and masking only during inference (Table VII). For both options, we used the mean value per feature from the training sets as the masking value for these experiments. When masking high perturb-ability features (Red/Red+Yellow) during both phases (Table VI), like Table V, performance remains nearly identical to the baseline across all models and datasets, for example, the ANN model on UNSW-NB15 retains 0.9879 accuracy and 0.9543 F1 score when masking Red+Yellow, matching the baseline’s 0.9883 accuracy and 0.9557 F1. This demonstrates that integrating masking into training preserves model efficacy without architectural changes. In contrast, inference-only masking (Table VII) shows a slight performance dip, as seen in the light blue cells, but maintains robustness ($>$ 0.94 F1) while enabling real-time adaptability. The minimal degradation in both scenarios validates that masking high-PS features, whether during training or dynamically at inference, effectively hardens models without compromising utility. Option B1’s training-integrated masking ensures stability, while B2’s inference-only approach prioritizes operational flexibility, allowing threat-responsive adjustments without retraining.
IX Mapping Traffic Morphing of problem-space attacks to ML-NIDS Features (RQ4)
To validate our proposed PS, we analyzed problem-space evasion adversarial attacks targeting ML-NIDS [11, 12, 13, 14, 15, 16, 17, 18]. Table VIII presents a detailed mapping of traffic morphing techniques employed in these problem-space adversarial attacks to the corresponding perturbed features within the feature-space of the two datasets utilized in this study. The table outlines the problem-space morphing techniques and maps them to the potentially impacted features. Additionally, it classifies these features according to our perturb-ability Score (PS) system, using a color-coded scheme: low perturb-ability (green), medium perturb-ability (yellow), and high perturb-ability (red).